Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on October 25th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • DC report
  • Magnet status
  • Decay solenoid status
  • Hydrogen status
  • Water status
    • Cooling water snapshots
    • Magnet current limits
  • Wednesday maintenance jobs 26th October
  • Shutdown activities : 28 October - 15 November (Colin)

Operations

  • MOM report and run plan to end of user cycle
    • Data-taking Summary to date: MiloradForKan.ods
    • Task list: Overall available run time : 60 hours (100k per 3 hours) from 6 am this morning to midnight on Thursday.
      Overall run time 60 hours
      Run time to date 9.5 hours
      Task Description Momentum # Triggers Triggers taken SSU (ECE) SSU M1/M2 FC SSD (ECE) SSD M1/M2 Beamline tag Cooling channel tag
      2 Scattering 140 300k at TOF2 111k at TOF2 Y Y/Y Solenoid Y N/N 3-140+M3-Test2 CCramp-6.1.4-NoTrims
      3 Scattering 170 300k at TOF2 106k at TOF2 Y Y/Y Solenoid Y N/N 3-170+M3-Test1 CCramp-6.1.4-NoTrims
      4 Scattering 200 300k at TOF2 ? Y Y/Y Solenoid Y N/N 3-200+M3-Test1 CCramp-6.1.4-NoTrims
      5 Scattering 240 300k at TOF2 ? Y Y/Y Solenoid Y N/N 3-240+M3-Test1 CCramp-6.1.4-NoTrims
  • Triggers taken with the empty absorber (no DS)
    Task Momentum # TOF2 Triggers
    2 140 26784
    3 170 35232
    4 200 30667
    5 240 38393
  • Overnight staffing
    • Tonight : Paul Kyberd / Steve Boyd
    • Wednesday night : Paul Kyberd / Ken Long (ending early at 7 am since Paul needs to leave then)
    • Thursday night : Mike Courthold to 6 am (then need someone to takeover) OR find two shifters for data-taking
  • Shutdown at 08:00 on Friday morning

Henry Report (probably in absentia)

Target: there were instances of locals twiddling laser powers without consulting the Target Expert (Joe), or even making a clear eLog entry. NOT ACCEPTABLE.

Recording of laser powers in eLog is intermittent... list of items to note is in the checklist, and if people aren't even bothering to follow the checklist...

Water:

Sheesh. I refer to slide 9 of my CM talk from barely 3 weeks ago:

  • where's the the eLog information requested in the shifters’ “what to do when ... fails” list?
  • why did the magnet power supplies actually trip?
  • was the situation in the trench checked against the most recent "magnet walkthrough" checklist?

From a quick conversation with John, I think we are getting glitches in the PLC, see also eLog 4044.

Why were the requested snapshots of the stripchart not being made? The striptool was obviously running, given that David did manage to include the tail-end of the recovery in his first upload. Let's say I'm disappointed at the way this has been recorded.

Just to be clear, I am not going to be bounced into some sort of constant remote oversight of MICE running to make sure that everyone is doing what they should be; not as Beamline Co-ordinator and not as DC.

I have provided procedures and limits to allow safe and reliable running of the equipment, and a few checklists to reinforce some procedures. That running happens from the MLCR, and it's for MICE to arrange for it to happen [the procedures and checklists to be followed by those in the control room, or else the equipment will not be made available]. _(Note added later: I drafted this offline, uploaded in a hurry and somehow lost the tail of the last paragraph. Now included it in brackets.)

Resulting To-do (in next run-up week):
  • Trench Water PLC - Control Glitches (Rob Searle)
  • Water Pump Control GUI - Expert Password (C&M)
  • Trench Waterleak detection
    • unroll sensor to full length to cover new plant (John G/Henry)
    • fix sensor readout(?); add to alarm handler (C&M)
  • Striptools for Target and Cooling Water
    • finalise templates (Henry)
    • add to Tiered app launcher (Paolo)

MiloradForKan.ods (25.5 KB) Boyd, Steven, 25 October 2016 14:31

OpsMeet25_10_2016.pdf - Tracker Summary M Uchida (255 KB) George, Melissa, 25 October 2016 14:45

Run8447imageData.pdf - 140 MeV, LiH, Channel on (2.31 MB) Rajaram, Durga, 25 October 2016 15:39

Run8449-tof-reco.pdf - TOF, 170 MeV (40.3 KB) Rajaram, Durga, 25 October 2016 15:48